# Magnetically-Driven Neutron-Rich Ejecta Unleashed: Global 3D Neutrino-GRMHD Simulations of Collapsars Reveal the Conditions for $r$-process Nucleosynthesis

### v1

@@ 4 Simulation results
Figure 5: Unbound mass binned by electron fraction in models with significant mass ejection (excluding pB0) at different times. Unbound mass is characterized according to the Bernoulli criterion in Eq. 18, where before jet launching we omit the magnetic field contribution ($b^{2}/\rho$ term). Stronger fields unbind more low $\,{Y_{\rm e}}$ material, maintaining a nearly flat distribution above the $\,{Y_{\rm e}}$ cutoff. Over time, the mixing of low $\,{Y_{\rm e}}$ outflows with the star raises the lower $\,{Y_{\rm e}}$ cutoff, resulting in an increased mass at higher $\,{Y_{\rm e}}$ values.
@@
Models pBs, pBw, which have lower mass accretion rates, produce smaller amounts of neutron-rich ejecta. Conversely, models with a flat density core (cBs, cBw, cBwA0) have higher accretion rates and can expel $\gtrsim 0.1\,\,{M_{\odot}}$ of roughly uniformly distributed mass in the range of $0.15\lesssim\,{Y_{\rm e}}\lesssim 0.5$ at $t\approx 0.1\,\,{{\rm s}}$. Stronger outflows (in model cBs) expel more neutron-rich ejecta while maintaining a similar flat $\,{Y_{\rm e}}$ distribution within the ejecta. The growth of the magnetic field in the disk halts the production of new low $\,{Y_{\rm e}}$ matter. At the same time, the previously generated neutron-rich ejecta mixes with the $\,{Y_{\rm e}}=0.5$ unshocked star, raising the $\,{Y_{\rm e}}$ in the ejecta. The two effects result in a continuous upward shift in the low $\,{Y_{\rm e}}$ distribution cutoff over time, increasing the mass content at the higher $\,{Y_{\rm e}}$ values. For example, at $t=0.1\,\,{{\rm s}}$, the uniform mass distribution of model cBs  has a cutoff at $\,{Y_{\rm e}}\approx 0.15$; by $t=0.1\,\,{{\rm s}}$, the cutoff has shifted to $\,{Y_{\rm e}}=0.25$.
@@
While this picture suggests that the ejecta will possess $\,{Y_{\rm e}}\approx 0.5$ by the time it reaches the stellar surface, there are several caveats about the mixing process. First, the low resolution at large distances from the disk leads to numerical diffusion and mixing that artificially increases the $\,{Y_{\rm e}}$ values. Higher resolution simulations are anticipated to show a milder evolution of the average $\,{Y_{\rm e}}$ in the ejecta. Second, the neutron rapid capture characteristic timescale of $t\lesssim 1\,\,{{\rm s}}$ (e.g., Lippuner & Roberts, 2017) indicates that hydrodynamic mixing over longer timescales will not alter the nuclear composition. The velocity of low $\,{Y_{\rm e}}$ outflows can reach a fraction of $c$, and the time dilation effect can effectively elongate the nucleosynthesis timescale. Models cBs, cBw, cBwA0 exhibit $M(\,{Y_{\rm e}}<0.3)\approx 1\,\,{M_{\odot}}$ at $t\approx 0.6\,\,{{\rm s}}$ after the low $\,{Y_{\rm e}}$ matter was ejected. This implies that the low $\,{Y_{\rm e}}$ ejecta has enough time to synthesize $r$-process nuclei, including lanthanides at $\,{Y_{\rm e}}\lesssim 0.25$ (e.g., Lippuner & Roberts, 2015), which will blend into the supernova ejecta, increasing the effective opacity and reddening the emission signature (e.g., Barnes & Metzger 2022; Barnes & Duffell 2023).
@@ 5 Summary
In this letter, we present the first global GRMHD simulations of collapsar explosions incorporating two-moment neutrino transport. We demonstrate that strongly magnetized disks eject massive ($M_{\rm ej}\approx 1\,\,{M_{\odot}}$) neutron-rich ($\,{Y_{\rm e}}<0.25$) outflows at extreme mass accretion rates ($\dot{M}\sim 10\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$). We also show that strong magnetic fields, necessary for outflow launching in collapsars, shorten the accretion timescales in the disk and thus suppress further disk neutronization.
@@
In our models, neutron-rich disk formation and ejection occur in a single episode. However, in nature, multiple cycles might take place if the magnetic flux in the envelope is more stochastic or if the angular momentum profile allows for the intermittent formation and destruction of the disk. Such repeated episodes could significantly increase the total amount of neutron-rich ejecta. Furthermore, if an intermittent jet progressively unbinds a larger fraction of the star, the neutron-rich ejecta could propagate outward with minimal mixing, similar to compact binary mergers, thereby enhancing the efficiency of heavy $r$-process nucleosynthesis. Alternatively, if the BH inherits its magnetic field from the proto-NS (Gottlieb et al., 2024), it could enable jet launching while maintaining moderate disk magnetization, potentially allowing for simultaneous disk neutronization and neutron-rich ejecta production. We plan to explore these scenarios in future work.
@@
In our simulations, we initiate exceptionally strong magnetic fields to resolve the MRI within the disk. This approach leads to high accretion speeds, forcing our models to maintain extremely high mass accretion rates of $\dot{M}\sim 10\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$ to facilitate disk neutronization. These high mass accretion rates also drive extreme jet energies of $10^{53}\,{\rm erg}\lesssim E_{\rm jet}\lesssim 10^{54}\,{\rm erg}$. While very atypical, the lower energy end of such extreme GRBs exist in nature (Burns et al., 2023; O’Connor et al., 2023), hinting at a possible connection of extremely bright GRBs with large neutron-rich ejecta mass. The high mass accretion rates will also cause the BH to double its mass within the collapse time, leading to a substantial spin-down (Lowell et al., 2024; Jacquemin-Ide et al., 2024), which complicates our assumption of static spacetime. Similarly, although our simulations produce up to $\,{M_{\rm ej}}\approx\,{M_{\odot}}$ of ejecta with $\,{Y_{\rm e}}<0.25$ within the freeze-out time, these results are influenced by the magnetic field strength and density profile of the collapsing star, which yield exceptionally high accretion rates and jet power in our models. Moreover, the final $\,{Y_{\rm e}}$ may change depending on what neutrino-matter interactions are taken into account in the neutrino transport, therefore the outcomes could be different if more sophisticated neutrino transport calculations are performed.
@@
In summary, our findings imply that only certain, special collapsar progenitors may contribute to the Galactic $r$-process enrichment. Hence, reduced estimates of $r$-process production from collapsars could better describe te observed Galactic abundances (see e.g., Rastinejad et al., 2024, for estimates). The precise mass and composition of neutron-rich ejecta across different progenitors remain uncertain. To address this, future studies must prioritize higher-resolution simulations that permit weaker magnetic fields, more moderate accretion rates and jet powers, allowing for the formation of neutron-rich ejecta under more realistic conditions. Stellar evolution models can provide such initial conditions for our models, which we aim to explore in the follow-up works. Furthermore, our current simulations lack crucial nucleosynthesis calculations with passive tracers to capture the thermodynamic evolution of the ejecta and do not include post-processing via nuclear reaction networks. These features are essential for determining whether the neutron-rich ejecta can undergo rapid neutron capture on seed nuclei before mixing with high-electron-fraction infalling material.
@@ Acknowledgements
We thank Irene Tamborra, Rodrigo Fernández, Philipp Mösta and Daniel Kasen for useful comments and discussions. DI is supported by Future Investigators in NASA Earth and Space Science and Technology (FINESST) award No. 80NSSC21K1851. OG acknowledges the Flatiron Research and CIERA Fellowships. B.D.M. acknowledges support from the National Science Foundation (grant number AST-2002577) and Simons Investigator Grant 727700. The Center for Computational Astrophysics at the Flatiron Institute is supported by the Simons Foundation. JJ acknowledges support by the NSF AST-2009884, NASA 80NSSC21K1746 and NASA XMM-Newton 80NSSC22K0799 grants. This work was performed in part at the Kavli Institute for Theoretical Physics (KITP) supported by grant NSF PHY-2309135. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC allocations m4603 (award NP-ERCAP0029085) and m2401. The computations in this work were, in part, run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation. An award of computer time was provided by the ASCR Leadership Computing Challenge (ALCC), Innovative and Novel Computational Impact on Theory and Experiment (INCITE), and OLCF Director’s Discretionary Allocation programs under award PHY129. AT acknowledges support by NASA 80NSSC22K0031, 80NSSC22K0799, 80NSSC18K0565 and 80NSSC21K1746 grants, and by the NSF grants AST-2009884, AST-2107839, AST-1815304, AST-1911080, AST-2206471, OAC-2031997.
Appendix A Neutrino transport implementation details
@@ A.1 Equations
We added the M1 neutrino transport scheme to the H-AMR  code (Liska et al., 2022) that integrates GRMHD equations of motion. In the presence of neutrino radiation, the evolution equations are:
@@
and $R^{\mu}_{\nu,s}$ is a stress-energy tensor of a single species of neutrinos. Here, we evolve 3 neutrino species $(\nu_{\rm e},\bar{\nu}_{\rm e},\nu_{\rm x})$, where the latter is the combination of 4 species $(\nu_{\mu},\bar{\nu}_{\mu},\nu_{\tau},\bar{\nu}_{\tau})$. We make this choice because the temperatures and neutrino energies in our collapsar simulations are too low and the formation of heavy lepton neutrinos is suppressed (e.g., Foucart et al., 2015). From here on, we imply the summation over all species and drop the index $s$.
In M1 method, we only consider the first two moments of the neutrino distribution function in so-called “gray” approximation, where we consider energy-integrated moments (see Shibata et al., 2011). Generally, the neutrino stress-energy tensor can be expressed through moments in the lab (coordinate) or fluid (comoving) rest frames:
@@
where $E,\,F^{\mu},\,P^{\mu\nu}$ are the energy, flux and stress tensor of the neutrino radiation measured by an observer in an inertial (zero angular momentum observer, or ZAMO) frame, $n^{\mu}$ is the 4-velocity of the ZAMO ($n_{\nu}=\left(-g^{tt}\right)^{-1/2}\delta^{t}_{\nu}$); similarly $J,\,H^{\mu},\,L^{\mu\nu}$ are the energy, flux and stress tensor of the neutrino radiation measured by an observer in the fluid rest frame, and $u^{\mu}$ is the fluid 4-velocity. Only the first two moments (energy and flux) are evolved, and to close the set of equations we need to express the stress tensor (A3) in terms of the first two moments (closure relation).
@@ A.2 Closure relation
In this work, we adopt a closure following Levermore (1984), where the main assumption is that there exists a “radiation” frame, an orthonormal frame in which the neutrino radiation is isotropic and satisfies the Eddington closure, which in the fluid frame gives:
@@
Following Sądowski et al. (2014); McKinney et al. (2014), we consider the decomposition of $R^{\mu}_{\nu}$ in terms of moments in so-called “radiation” frame:
@@
This closure reproduces the optically thin and thick limits relatively well. However, unsurprisingly, it breaks down in the presence of colliding beams because of the key assumption that a single radiation frame exists. Since only one direction - that of the boost - is distinguished, in case of intersecting beams, the scheme approximates the radiation having the mean direction (Sądowski et al., 2014). Alternative analytic closures suffer from similar problems (see Murchikova et al., 2017, for an overview).
@@ A.3 Neutrino evolution equations
We rewrite the evolution equation A1 as:
@@
where $\eta,\kappa_{\rm a},\kappa_{\rm s}$ are gray (energy-integrated) emissivities, absorption and scattering opacities, described in A.4.4. Expressed in conservative form, Eq. A6 becomes:
@@
In addition, we also evolve the neutrino number density to accurately track the composition evolution following the approach proposed by Foucart et al. (2016); Radice et al. (2022). The neutrino number density evolution equation is:
@@
where $\hat{N}$ is the neutrino number density in the fluid frame $\hat{N}=-N^{\mu}u_{\mu}$, $\eta_{\rm N},\kappa_{\rm N}$ are the neutrino number emissivity and absorption coefficients (see A.4.4). We adopted the form:
@@
for the neutrino number current under the assumption that the neutrino number and energy flux are aligned which is exact if neutrinos had a single energy (Radice et al., 2022). The composition evolution is then (from lepton number conservation):

### v2

@@ 4 Simulation results
Figure 5: Unbound mass binned by electron fraction in models with significant mass ejection (excluding pB0) at different times. Unbound mass is characterized according to the Bernoulli criterion in Eq. 18, where before jet launching, we omit the magnetic field contribution ($b^{2}/\rho$ term). Stronger fields unbind more low $\,{Y_{\rm e}}$ material, maintaining a nearly flat distribution above the $\,{Y_{\rm e}}$ cutoff. Over time, the mixing of low $\,{Y_{\rm e}}$ outflows with the star raises the lower $\,{Y_{\rm e}}$ cutoff, resulting in an increased mass at higher $\,{Y_{\rm e}}$ values.
@@
Models pBsand pBw, which have lower mass accretion rates, produce smaller amounts of neutron-rich ejecta. Conversely, models with a flat density core (cBs, cBw, and cBwA0) have higher accretion rates and can expel $\gtrsim 0.1\,\,{M_{\odot}}$ of roughly uniformly distributed mass in the range of $0.15\lesssim\,{Y_{\rm e}}\lesssim 0.5$ at $t\approx 0.1\,\,{{\rm s}}$. Stronger outflows (in model cBs) expel more neutron-rich ejecta while maintaining a similar flat $\,{Y_{\rm e}}$ distribution within the ejecta. The growth of the magnetic field in the disk halts the production of new low $\,{Y_{\rm e}}$ matter. At the same time, the previously generated neutron-rich ejecta mixes with the $\,{Y_{\rm e}}=0.5$ unshocked star, raising the $\,{Y_{\rm e}}$ in the ejecta. The two effects result in a continuous upward shift in the low $\,{Y_{\rm e}}$ distribution cutoff over time, increasing the mass content at the higher $\,{Y_{\rm e}}$ values. For example, at $t=0.1\,\,{{\rm s}}$, the uniform mass distribution of model cBs  has a cutoff at $\,{Y_{\rm e}}\approx 0.15$; by $t=0.1\,\,{{\rm s}}$, the cutoff has shifted to $\,{Y_{\rm e}}=0.25$.
@@
While this picture suggests that the ejecta will possess $\,{Y_{\rm e}}\approx 0.5$ by the time it reaches the stellar surface, there are several caveats about the mixing process. First, the low resolution at large distances from the disk leads to numerical diffusion and mixing that artificially increases the $\,{Y_{\rm e}}$ values. Higher-resolution simulations are anticipated to show a milder evolution of the average $\,{Y_{\rm e}}$ in the ejecta. Second, the neutron rapid capture characteristic timescale of $t\lesssim 1\,\,{{\rm s}}$ (e.g., Lippuner & Roberts, 2017) indicates that hydrodynamic mixing over longer timescales will not alter the nuclear composition. The velocity of low $\,{Y_{\rm e}}$ outflows can reach a fraction of $c$, and the time dilation effect can effectively elongate the nucleosynthesis timescale. Models cBs, cBw, and cBwA0 exhibit $M(\,{Y_{\rm e}}<0.3)\approx 1\,\,{M_{\odot}}$ at $t\approx 0.6\,\,{{\rm s}}$ after the low $\,{Y_{\rm e}}$ matter was ejected. This implies that the low $\,{Y_{\rm e}}$ ejecta has enough time to synthesize $r$-process nuclei, including lanthanides at $\,{Y_{\rm e}}\lesssim 0.25$ (e.g., Lippuner & Roberts, 2015), which will blend into the supernova ejecta, increasing the effective opacity and reddening the emission signature (e.g., Barnes & Metzger 2022; Barnes & Duffell 2023).
@@ 5 Summary
In this Letter, we present the first global GRMHD simulations of collapsar explosions incorporating M1 neutrino transport. We demonstrate that strongly magnetized disks eject massive ($M_{\rm ej}\approx 1\,\,{M_{\odot}}$), neutron-rich ($\,{Y_{\rm e}}<0.25$) outflows at extreme mass accretion rates ($\dot{M}\sim 10\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$). We also show that strong magnetic fields, necessary for outflow launching in collapsars, shorten the accretion timescales in the disk, simultaneously growing the neutrino-cooling timescales due to decreasing densities and thus suppress further disk neutronization.
@@
In our models, neutron-rich disk formation and ejection occur in a single episode. However, in nature, multiple cycles might take place if the magnetic flux in the envelope is more stochastic, or if the angular momentum profile allows for the intermittent formation and destruction of the disk. Such repeated episodes could significantly increase the total amount of neutron-rich ejecta. Furthermore, if an intermittent jet progressively unbinds a larger fraction of the star, the neutron-rich ejecta could propagate outward with minimal mixing, similar to compact binary mergers, thereby enhancing the efficiency of heavy $r$-process nucleosynthesis. Alternatively, if the BH inherits its magnetic field from the proto-NS (Gottlieb et al., 2024), it could enable jet launching while maintaining moderate disk magnetization, potentially allowing for simultaneous disk neutronization and neutron-rich ejecta production. We plan to explore these scenarios in future work.
@@
In our simulations, we initiate exceptionally strong magnetic fields to resolve the MRI within the disk. This approach leads to high disk inflow velocities, forcing our models to maintain extremely high mass accretion rates of $\dot{M}\sim 10\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$ to facilitate disk neutronization. These high mass accretion rates also drive extreme jet energies of $10^{53}\,{\rm erg}\lesssim E_{\rm jet}\lesssim 10^{54}\,{\rm erg}$. While very atypical, the lower-energy end of such extreme GRBs exists in nature (Burns et al., 2023; O’Connor et al., 2023), hinting at a possible connection of extremely bright GRBs with large neutron-rich ejecta mass. The high mass accretion rates will also cause the BH to double its mass within the collapse time, leading to a substantial spin-down (Lowell et al., 2024; Jacquemin-Ide et al., 2024), which complicates our assumption of static spacetime. Similarly, although our simulations produce up to $\,{M_{\rm ej}}\approx\,{M_{\odot}}$ of ejecta with $\,{Y_{\rm e}}<0.25$ within the freeze-out time, these results are influenced by the magnetic field strength and density profile of the collapsing star, which yield exceptionally high accretion rates and jet power in our models. Moreover, the final $\,{Y_{\rm e}}$ may change depending on what neutrino-matter interactions are taken into account in the neutrino transport; therefore, the outcomes could be different if more sophisticated neutrino-transport calculations are performed.
@@
In summary, our findings imply that only certain special collapsar progenitors may contribute to the Galactic $r$-process enrichment. Hence, reduced estimates of $r$-process production from collapsars could better describe the observed Galactic abundances (see e.g., Rastinejad et al., 2024, for estimates). The precise mass and composition of neutron-rich ejecta across different progenitors remain uncertain. To address this, future studies must prioritize higher-resolution simulations that permit weaker magnetic fields, more moderate accretion rates, and jet powers, allowing for the formation of neutron-rich ejecta under more realistic conditions. Stellar evolution models can provide such initial conditions for our models, which we aim to explore in the follow-up works. Furthermore, our current simulations lack crucial nucleosynthesis calculations with passive tracers to capture the thermodynamic evolution of the ejecta and do not include postprocessing via nuclear reaction networks. These features are essential for determining whether the neutron-rich ejecta can undergo $r$-process on seed nuclei before mixing with high-electron-fraction infalling material.
@@ Acknowledgements
We thank Irene Tamborra, Rodrigo Fernández, Philipp Mösta, and Daniel Kasen for useful comments and discussions. We thank the anonymous reviewer for their feedback that led to improvements throughout the manuscript. D.I. thanks Aretaios Lalakos and Nick Kaaz for fruitful discussions. D.I. is supported by Future Investigators in NASA Earth and Space Science and Technology (FINESST) award No. 80NSSC21K1851. O.G. acknowledges the Flatiron Research and CIERA Fellowships. B.D.M. acknowledges support from the National Science Foundation (grant No. AST-2002577) and Simons Investigator grant 727700. The Center for Computational Astrophysics at the Flatiron Institute is supported by the Simons Foundation. J.J. acknowledges support by the NSF AST-2009884, NASA 80NSSC21K1746 and NASA XMM-Newton 80NSSC22K0799 grants. This work was performed in part at the Kavli Institute for Theoretical Physics (KITP) supported by grant NSF PHY-2309135. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC allocations m4603 (award NP-ERCAP0029085) and m2401. The computations in this work were, in part, run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation. An award of computer time was provided by the ASCR Leadership Computing Challenge (ALCC), Innovative and Novel Computational Impact on Theory and Experiment (INCITE), and OLCF Director’s Discretionary Allocation programs under award PHY129. A.T. acknowledges support by NASA 80NSSC22K0031, 80NSSC22K0799, 80NSSC18K0565 and 80NSSC21K1746 grants, and by the NSF grants AST-2009884, AST-2107839, AST-1815304, AST-1911080, AST-2206471, AST-2407475, OAC-2031997.
Appendix A Neutrino-transport implementation details
@@ A.1 Equations
We added the M1 neutrino-transport scheme to the H-AMR  code (Liska et al., 2022) that integrates GRMHD equations of motion. In the presence of neutrino radiation, the evolution equations are
@@
and $R^{\mu}_{\nu,s}$ is a stress-energy tensor of a single species of neutrinos. Here, we evolve three neutrino species $(\nu_{\rm e},\bar{\nu}_{\rm e},\nu_{\rm x})$, where the latter is the combination of four species $(\nu_{\mu},\bar{\nu}_{\mu},\nu_{\tau},\bar{\nu}_{\tau})$. We make this choice because the temperatures and neutrino energies in our collapsar simulations are too low and the formation of heavy lepton neutrinos is suppressed (e.g., Foucart et al., 2015). From here on, we imply the summation over all species and drop the index $s$.
In M1 method, we only consider the first two moments of the neutrino distribution function in the so-called “gray” approximation, where we consider energy-integrated moments (see Shibata et al., 2011). Generally, the neutrino stress-energy tensor can be expressed through moments in the lab (coordinate) or fluid (comoving) rest frames:
@@
where $E,\,F^{\mu},\,andP^{\mu\nu}$ are the energy, flux, and stress tensor of the neutrino radiation measured by an observer in an inertial (zero angular momentum observer, or ZAMO) frame and $n^{\mu}$ is the 4-velocity of the ZAMO ($n_{\nu}=\left(-g^{tt}\right)^{-1/2}\delta^{t}_{\nu}$); similarly $J,\,H^{\mu},\,andL^{\mu\nu}$ are the energy, flux, and stress tensor of the neutrino radiation measured by an observer in the fluid rest frame, and $u^{\mu}$ is the fluid 4-velocity. Only the first two moments (energy and flux) are evolved, and to close the set of equations, we need to express the stress tensor (Eq. (A3)) in terms of the first two moments (closure relation).
@@ A.2 Closure relation
In this work, we adopt a closure following Levermore (1984), where the main assumption is that there exists a “radiation” frame, an orthonormal frame in which the neutrino radiation is isotropic and satisfies the Eddington closure, which in the fluid frame gives
@@
Following Sądowski et al. (2014) and McKinney et al. (2014), we consider the decomposition of $R^{\mu}_{\nu}$ in terms of moments in the so-called “radiation” frame:
@@
This closure reproduces the optically thin and thick limits relatively well. However, unsurprisingly, it breaks down in the presence of colliding beams because of the key assumption that a single radiation frame exists. Since only one direction – that of the boost – is distinguished, in case of intersecting beams, the scheme approximates the radiation having the mean direction (Sądowski et al., 2014). Alternative analytic closures suffer from similar problems (see Murchikova et al., 2017, for an overview).
@@ A.3 Neutrino evolution equations
We rewrite the evolution equation A1 as
@@
where $\eta,\kappa_{\rm a},and\kappa_{\rm s}$ are gray (energy-integrated) emissivities, absorption, and scattering opacities, described in A.4.4. Expressed in conservative form, Eq. A6 becomes
@@
In addition, we also evolve the neutrino number density to accurately track the composition evolution following the approach proposed by Foucart et al. (2016) and Radice et al. (2022). The neutrino number density evolution equation is
@@
where $\hat{N}$ is the neutrino number density in the fluid frame $\hat{N}=-N^{\mu}u_{\mu}$, $\eta_{\rm N},\kappa_{\rm N}$ are the neutrino number emissivity and absorption coefficients (see A.4.4). We adopted the form
@@
for the neutrino number current under the assumption that the neutrino number and energy flux are aligned, which is exact if neutrinos had a single energy (Radice et al., 2022). The composition evolution is then (from lepton number conservation)